OpenStudy (anonymous):

10.04] What is the exact value of the expression the square root 72. − the square root of 8. + the square root of 128.? Simplify if possible. 8the square root of 2. 12the square root of 2. 8the square root of 3. 12the square root of 3. [10.04] What is the exact value of the expression 6the square root of y to the eighth power. − the square root of 9 times y to the fourth power. + 3the square root of y to the eighth power. − the square root of y to the fourth power.? Simplify if possible. y4 9y4 − 4y2 3y4 − 10y2 y2

OpenStudy (dumbcow):

this hurts my eyes...use math symbols, dont write everything out

OpenStudy (anonymous):

\[\sqrt{72}-\sqrt{8}+\sqrt{128}=6 \sqrt{2}-2 \sqrt{2}+8 \sqrt{2}=12 \sqrt{2} \]

OpenStudy (dumbcow):

basically you are simplifying radicals by pulling out perfect squares use propertys: \[\sqrt{m*n} = \sqrt{m} \sqrt{n}\] \[\sqrt{m^{2}} = m\]

OpenStudy (anonymous):

so whats the answer